OpenStudy (shesolitt):

I WILL MEDAL AND FAN!! PLS HELP Write the equation of the line that passes through (3, 4) and (2, –1) in slope-intercept form. y = 3x – 7 y = 3x – 5 y = 5x – 11 y = 5x – 9

OpenStudy (shesolitt):

@likeabossssssss

OpenStudy (shesolitt):

@FaithDaLibra @mathmath333 @allykins

OpenStudy (anonymous):

\[\text{Solve}\left[\frac{y-4}{x-3}=\frac{-5}{-1},y\right] \]y = 5x-11